# Signature Isolation Forest

### v3

@@ 4.1 Parameters Sensitivity Analysis
Methods : SIF K-SIFGW FIFGW K-SIFC FIFC K-SIFB FIFB IF OCSVM fHD fSDO
Chinatown 1 0.90 0.84 0.99 0.9 1 0.83 0.81 0.6 0.73 1
Coffee 0.84 0.92 0.87 0.85 0.91 0.83 0.60 0.74 0.76 0.89 0.9
ECGFiveDays 0.93 0.90 0.88 0.92 0.95 0.90 0.93 0.85 0.93 0.84 0.96
ECG200 0.85 0.82 0.85 0.85 0.86 0.83 0.82 0.85 0.81 0.83 0.89
Handoutlines 0.84 0.83 0.83 0.83 0.83 0.82 0.8 0.84 0.83 0.82 0.83
SonyRobotAI1 0.99 0.96 0.98 0.95 0.93 0.95 0.76 0.96 0.96 0.88 0.96
SonyRobotAI2 0.93 0.89 0.88 0.92 0.80 0.93 0.84 0.87 0.84 0.84 0.88
StarLightCurves 0.80 0.75 0.71 0.76 0.67 0.76 0.77 0.72 0.71 0.71 0.79
TwoLeadECG 0.92 0.92 0.76 0.92 0.74 0.92 1 0.78 0.54 0.63 1
ECG5000 0.90 0.92 0.89 0.97 0.88 0.91 0.88 0.94 0.94 0.97 0.92
@@ 4.3 Real-data Anomaly Detection Benchmark
K-SIF is examined in the context of three introduced dictionaries: Brownian, Cosine and Gaussian wavelets, as done for FIF. The parameters are configured with $N=100$, $m=\text{min}(256,n)$, the height limit set to $\lceil\log_{2}(m)\rceil$ for both FIF and K-SIF/SIF. The number of split windows is fixed with $\omega=10$, and the depth is fixed with $k=3$. We compare the K-SIF/SIF methods against two widely used multivariate anomaly detection techniques and two functional depths with default settings. The multivariate methods, namely, Isolation Forest (IF; Liu et al.,, 2008) and one-class support vector machine (OCSVM; Schölkopf et al.,, 2001), are applied following dimension reduction via Functional PCA. We retain 20 principal components with the largest eigenvalues using the Haar basis. The depths considered include the random projection halfspace depth (Cuevas et al.,, 2007) and the functional Stahel-Donoho outlyingness (Hubert et al.,, 2015). Due to the relative small size of datasets, see Table 1 in the Appendix, deep learning based method performs poorly as shown in Tables 5, 6 and 7 in the Appendix.
@@ 5 Discussion & Conclusion
Acknowledgments
Guillaume Staerman thanks the G-research grant for funding this work.

### v4

@@ 3 Signature Isolation Forest Methods
Figure 1: Geometric visualization of depth-2 signature terms, where $S^{(1,2)}$ (cyan region) and $S^{(2,1)}$ (purple region) represent areas corresponding to coordinate signatures. The displacement term $\Delta X_{1}$ and $\Delta X_{2}$ along each axis capture the depth-1 terms of the transform.
@@ 4.1 Parameters Sensitivity Analysis
Methods: SIF K-SIFGW K-SIFC K-SIFB IF OCSVM fHD fSDO DeepSVDD AnoGan AutoEncoder VAE
Chinatown 1 0.90 0.99 1 0.81 0.6 0.73 1 0.5 0.72 1 0.99
Coffee 0.84 0.92 0.85 0.83 0.74 0.76 0.89 0.9 0.55 0.5 0.83 0.83
ECGFiveDays 0.93 0.90 0.92 0.90 0.85 0.93 0.84 0.96 0.67 0.79 0.96 0.96
ECG200 0.85 0.82 0.85 0.83 0.85 0.81 0.83 0.89 0.5 0.72 0.87 0.86
Handoutlines 0.84 0.83 0.83 0.82 0.84 0.83 0.82 0.83 0.63 0.55 0.82 0.82
SonyRobotAI1 0.99 0.96 0.95 0.95 0.96 0.96 0.88 0.96 0.52 0.91 0.95 0.95
SonyRobotAI2 0.93 0.89 0.92 0.93 0.87 0.84 0.84 0.88 0.64 0.87 0.88 0.88
StarLightCurves 0.80 0.75 0.76 0.76 0.72 0.71 0.71 0.79 0.60 0.30 0.76 0.77
TwoLeadECG 0.92 0.92 0.92 0.92 0.78 0.54 0.63 1 0.46 0.68 1 1
ECG5000 0.90 0.92 0.97 0.91 0.94 0.94 0.97 0.92 0.59 0.65 0.92 0.92
Methods: SIF KSIF FIF IF OCSVM fHD fSDO DeepSVDD AnoGan AutoEncoder VAE
Chinatown 0.3 0.5 1.3 0.3 0.3 0.1 0.1 18 78 91 78
Coffee 0.8 1 1.6 0.3 0.3 0.1 0.1 16 88 91 90
ECGFiveDays 0.6 0.6 1.2 0.5 0.6 0.3 0.3 21 84 87 89
ECG200 3 5 9 1 1 0.7 0.7 18 151 91 90
Handoutlines 30 50 90 7 7 2 2 27 681 593 365
SonyRobotAI1 0.8 0.8 1.6 0.5 0.6 0.3 0.3 18 84 95 84
SonyRobotAI2 0.9 0.8 1.6 0.5 0.6 0.3 0.3 19 86 92 82
StarLightCurves 13 20 62 3 5 1 1 22 500 202 175
TwoLeadECG 0.6 0.6 1 0.3 0.3 0.2 0.2 9 75 66 58
ECG5000 5.4 4 15 1 1 0.6 0.6 9 270 102 91
Table 2: Computational time in seconds of different anomaly detection methods calculated on the test set.
@@ 4.3 Real-data Anomaly Detection Benchmark
K-SIF is examined in the context of three introduced dictionaries: Brownian, Cosine and Gaussian wavelets, as done for FIF. The parameters are configured with $N=100$, $m=\text{min}(256,n)$, the height limit set to $\lceil\log_{2}(m)\rceil$ for both FIF and K-SIF/SIF. The number of split windows is fixed with $\omega=10$, and the depth is fixed with $k=3$. We compare the K-SIF/SIF methods against two widely used multivariate anomaly detection techniques and two functional depths with default settings. The multivariate methods, namely, Isolation Forest (IF; Liu et al.,, 2008) and one-class support vector machine (OCSVM; Schölkopf et al.,, 2001), are applied following dimension reduction via Functional PCA. We retain 20 principal components with the largest eigenvalues using the Haar basis. The depths considered include the random projection halfspace depth (Cuevas et al.,, 2007) and the functional Stahel-Donoho outlyingness (Hubert et al.,, 2015).
@@
Our proposed methods, SIF and KSIF, demonstrate strong computational efficiency compared to both traditional and deep learning-based anomaly detection methods. As shown in Table 6, SIF and KSIF consistently achieve lower execution times than deep learning approaches, such as AutoEncoder and AnoGan, which require significantly more computational resources. Notably, on datasets like HandOutlines and StarLightCurves, KSIF and SIF exhibit a substantial reduction in processing time compared to deep learning models, making them more suitable for real-time or resource-constrained applications. While traditional methods like IF and OCSVM offer comparable efficiency, our methods maintain a favorable balance between speed and detection performance. These results highlight the advantage of our approach in scenarios where rapid anomaly detection is critical, reinforcing its practicality for real-world deployment.